OpenStudy (anonymous):

Please help, how do i get the derivative of f(x)= x^(1/3). I started but got stuck

OpenStudy (anonymous):

\[\frac{ d x ^{1/3} }{ dx }\]

OpenStudy (anonymous):

f'(x)= 1/3 (x) ^1/3-1 =1/3(x)^-2/3

OpenStudy (anonymous):

\[\frac{ d x^n }{d x } = n x ^{n-1}\]

OpenStudy (anonymous):